OpenStudy (anonymous):

WILL GIVE A MEDAL The hypotenuse of a right triangle is 7 inches, and one of the legs is sqrt 13 inches. Find the length of the other leg. A. 6 inches B. sqrt 62 inches C. sqrt 128 inches

OpenStudy (anonymous):

From pythagoras theorem: we have (length of hypotenuse)^2 - (leg 1)^2 = (leg 2)^2. So we have 7^2 - (sqrt13)^2= (leg 2)^2

OpenStudy (anonymous):

@shawmoes So... what do I do from here? I get really confused when they throw squareroots in the problem. :/

OpenStudy (anonymous):

When you square a number with a square root, you just get that number underneath the square root. So

OpenStudy (anonymous):

Oh, okay! So... 7^2 - 13 = c?

OpenStudy (anonymous):

7^2 - 13 = c^2

OpenStudy (anonymous):

#sayNOtoPERFUMESotherTHANattar

OpenStudy (anonymous):

c^2= 36?? Is that correct?

OpenStudy (anonymous):

Yes !

OpenStudy (anonymous):

So what is c ?

OpenStudy (anonymous):

So it be 6?

OpenStudy (anonymous):

Correct :) . Good job.

OpenStudy (anonymous):

Oh, thank you so much!!